OpenStudy (anonymous):

Which of the following correctly completes the square for the equation below? x2 - 9x = 0

OpenStudy (anonymous):

If you have x^2+18x=0 then you can write it as x^2+18x+81=81 since I've added 81 to both sides the equation is still complete. With x^2+18x+81 you can factor that down to (x+9)^2 since 9*2=18 and 9^2=81. Hence the whole thing is (x+9)^2=81 Dean M · 5

OpenStudy (anonymous):

im confused

MsBrains (ms-brains):

What are you confused about?

OpenStudy (anonymous):

the whole thing

OpenStudy (ijlal):